Evaluate the integral of (2+e^y)/(e^y) dy


1
Expert's answer
2021-02-24T11:09:17-0500

"\begin{array}{l}
\int {\frac{{2 + {e^y}}}{{{e^y}}}dy = } \int {\left( {\frac{2}{{{e^y}}} + \frac{{{e^y}}}{{{e^y}}}} \right)} dy = \int {2{e^{ - y}}dy + \int {1dy = } } - \int {2{e^{ - y}}d( - y) + \int {1dy = } } - 2{e^{ - y}} + y + C = \\
 = - \frac{2}{{{e^y}}} + y + C
\end{array}"